# CLARA: A Modular Framework for Unsupervised Transit Detection Using TESS Light Curves

### v1

@@ 4.2 Weighted Root Sum Score: A revision of URF anomaly scores
To give higher preference to better scoring, we set the weights as $w_{3}=0.8$ and $w_{9}=0.2$. For a standardized comparison across urf-4 model runs on different test sets, we normalize $s_{w}$ (N-WRSS) using:
@@
Figure 15: Distribution of N-WRSS g.t.e 0.4 on the intersection set of $\alpha=0.3$ and $\alpha=0.9$ models run on TESS SPOC Sector 1 Curves
Figure 16: Distribution of N-WRSS g.t.e 0.6 on the intersection set of $\alpha=0.3$ and $\alpha=0.9$ models run on TESS SPOC Sector 1-5 Curves
@@
The normalized weighted root sum of squares (N-WRSS $\hat{s}_{w}$) provides a joint metric that captures the agreement or divergence between different URF models ($\alpha_{0.3}$ and $\alpha_{0.9}$ in this case). By computing $\hat{s}_{w}$ over the intersection set - i.e., light curves flagged by both models - we:
@@
  • Suppress model-specific bias: Since $\alpha_{0.3}$ and $\alpha_{0.9}$ may differ in sensitivity due to their hyperparameter tuning or training noise characteristics, $\hat{s}_{w}$ acts as a joint view, minimizing the influence of any one model’s noise.
  • Support comparative benchmarking: It gives a unified score to evaluate intersection anomalies, useful when testing the consistency or precision of later similarity-based matching (e.g., cosine similarity labeling).
Thus, $\hat{s}_{w}$ on the $\alpha_{0.3}$ - $\alpha_{0.9}$ intersection forms a critical bridge for robust scoring before downstream similarity analysis or labeling.
In its current implementation, the N-WRSS reduces score variance and compresses outlier influence in a computationally tractable way. For further mathematical and theoretical justification of this, refer to the Theoretical Framework Development section of Section 6.5.
@@ 4.3 Astrometric and Stellar Parameter Augmentation
To enable physical interpretation of anomalies detected by URF-4, we enrich high-scoring candidates with astrophysical and astrometric metadata. Specifically, we extract relevant features from the TESS Input Catalog (TIC) and Gaia DR3 for all candidates exceeding a chosen anomaly threshold (in our case N-WRSS >=>= 0.4).
@@
Refer to Appendix B for a descriptive list of these features.
@@ 4.4 Evaluating Astrometric Feature Discriminability
To assess which combinations of astrometric and stellar parameters most effectively distinguish high-anomaly candidates, we perform a t-SNE embedding followed by silhouette score ranking of feature sets on Sector 1 URF-4 $\alpha_{0.3}\cap\alpha_{0.9}$ set.
@@
The silhouette score provides a measure of cluster separation, with higher values indicating that the selected feature combination better separates anomalous from typical light curves in the reduced space.
This process identifies the most discriminative physical features for downstream interpretation, linking anomaly detection outcomes to astrophysical structure.
These findings provide an initial quantitative basis for linking astrometric and stellar parameters to anomaly clustering behavior. Future work (6.5) will expand this into a formal statistical validation framework, including comparison with established methods (e.g. Crake & Martínez-Galarza (2023); Melton et al. (2024)) and deeper analysis of why certain features like Teff, RUWE, and $v_{\mathrm{tan}}$ exhibit strong correlations with high N-WRSS anomalies, thereby enabling more grounded astrophysical interpretation of detected outliers.
4.5 Results
  • The top-performing set was: [’pmdec’, ’Teff’, ’vtan’, ’ruwe’] with a silhouette score of  0.069. Top 10 feature sets are provided in Table 6
  • The most visually separable and physically interpretable set was: [’Teff’, ’vtan’, ’ruwe’], showing excellent concentrated red region clusters.
  • Simpler sets such as [’Teff’, ’ruwe’] also performed well, confirming the significance of these features.
Table 6: Top 10 astrometric feature sets ranked by silhouette score based on their separation of high- and low-anomaly candidates in t-SNE space.
Rank Feature Set Silhouette Score
1 [pmdec, Teff, vtan, ruwe] 0.069
2 [Teff, ruwe] 0.068
3 [pmra, pmdec, Teff] 0.064
4 [pmra, rad] 0.062
5 [rad, vtan] 0.056
6 [Teff, rad] 0.055
7 [plx, pmdec, rad, logg] 0.053
8 [pmra, pmdec, rad, logg] 0.052
9 [plx, pmra, rad, lum] 0.051
10 [pmra, pmdec, Teff, vtan, ruwe] 0.051
Figure 17: t-SNE clustering of Sector 1 N-WRSS >= 0.6 curves based on [pmdec, Teff, vtan, ruwe]. N-WRSS goes from blue to red increasing from 0.0 to 0.9 (non-linearly to mark N-WRSS >= 0.6 better visually)
Figure 18: t-SNE clustering of Sector 1 N-WRSS >= 0.6 curves based on [Teff, vtan, ruwe]. N-WRSS goes from blue to red increasing from 0.0 to 0.9 (non-linearly to mark N-WRSS >= 0.6 better visually)
4.6 Next Steps: Cross-Sector Consistency and Physical Interpretation
To build on the silhouette-based feature evaluation, we propose a focused cross-sector analysis using two key feature sets: the top-scoring set [pmdec, Teff, vtan, ruwe] and the most visually separable set [Teff, vtan, ruwe]. We will apply these feature combinations to high-N-WRSS URF-4 $\alpha_{0.3}\cap\alpha_{0.9}$ candidates across four TESS sectors to evaluate the consistency of anomaly clustering in astrometric and stellar parameter space.
The objective is to determine whether URF-identified high scoring anomalies consistently occupy specific, distinguishable regions in this astrophysical feature space.
This multi-sector analysis will serve as an important validation step, linking unsupervised anomaly detection results to underlying physical features and assessing the reproducibility of clustering structures in high-dimensional parameter space. Results of this cross sector run are presented in figures 19 and 20.
4.7 Interpretation of Cross-Sector Embeddings
The t-SNE embeddings in Figures 19 and 20 reveal consistent structural patterns across TESS Sectors 1–4. In particular, the [Teff, vtan, ruwe] projection shows a recurring concentration of high N-WRSS scores (red regions) along narrow manifolds in the reduced space, suggesting the existence of a physically distinct anomaly subpopulation. The inclusion of proper motion in the extended feature set [pmdec, Teff, vtan, ruwe] enhances this separability, with clearer filamentary structures observed. This reinforces the hypothesis that high-scoring anomalies are not randomly distributed but instead correlate with specific stellar or kinematic regimes, validating the physical relevance of the URF-derived anomaly scores across sectors.
@@ 4.8 SIMBAD Weak Supervision and Morphological Similarity Matching
To bridge anomaly detection with astrophysical classification, we introduce a weakly supervised labeling approach using the SIMBAD object taxonomy. A subset of high-anomaly light curves (TOIs from Sector 2 - specific choice of sector in this case is arbitrary - for which Simbad Object Labels could be found) is cross-matched with MAST astronomical coordinates to retrieve SIMBAD labels such as Pl, EB*, V*, and others. This offers valuable physical priors without introducing strong biases. These labeled instances are used to construct a Sector 2 reference set of TOIs with known SIMBAD classifications, enabling downstream morphological matching and interpretation.
We extract feature vectors from these labeled TOIs using Lomb-Scargle periodograms and flux values of the curves as described in 3.1, apply PCA for dimensionality reduction, and save the compressed representations. These features form a labeled reference basis for similarity-based classification.
Subsequently, we compute pairwise cosine similarity between urf-4 $\alpha_{0.3}\cap\alpha_{0.9}$ set candidates and this labeled reference set. Each anomaly is assigned the closest-matching SIMBAD label group based on morphological similarity. Following label groups have been defined for this purpose:
@@ 4.8.2 Morphological Classification via Cosine Similarity
    Extracting feature vectors from the input light curve using the same pipeline used for the labeled TOIs as discussed in Section 3.1.
@@
This classification approach allows for the astrophysically guided annotation of anomaly candidates, making it possible to investigate the clustering of morphologically similar signals and to validate the physical realism of urf-identified anomalies.
4.8.3 Results of Morphological Classification via Cosine Similarity
To assess the astrophysical plausibility of morphologically matched labels, we analyze the distribution of URF anomaly scores across the assigned SIMBAD label categories.
  1. 1.
  2. 2.
  3. 3.
4.8.4 Interpretation of Cosine Similarity Results
The three diagnostic views (figures 21, 22 and 23) enable a structured assessment of how anomaly scores and morphological similarity align with known astrophysical classes.
Second, the high-similarity/low-anomaly set (Figure 22) demonstrates that URF correctly assigns low anomaly scores to morphologically standard light curves which can be classified into a labeled category. This supports the argument that low URF scores represent "normal" behavior in the training distribution.
4.8.5 Cosine Similarity Label Group Statistics
To evaluate the distribution of anomaly strength across label groups, we stratify all matched candidates by their assigned label group (planet_like, stellar, binary_star) and analyze how the number of candidates varies with increasing N-WRSS (normalized weighted root sum score) threshold. The plots in figures 24, 25 and 26, for each label group, show the number of candidates with anomaly score above a given threshold.
Figure 24: Planet-like candidates above n-wrss threshold for sectors 1 to 5
Figure 25: Stellar candidates above n-wrss threshold for sectors 1 to 5
Figure 26: Binary-star candidates above n-wrss threshold for sectors 1 to 5
Figure 27: First curve in cluster 1 of dbscan clustering on dtw matrix with eps 4.5 and min_samples 2
Figure 28: Second curve in cluster 1 of dbscan clustering on dtw matrix with eps 4.5 and min_samples 2
Figure 29: Silhouette score heatmap for grid search over eps and min_sample values. Two optimal values are eps = 2.0, min_samples = 10 and eps = 4.5, min_samples = 2
@@ 4.9 Clustering and Morphological Filtering Methods
To identify transit-like morphologies, we selected the planet_like classified group of curves from TESS Sectors 1 through 5. We applied a threshold on the normalized weighted root sum square (N-WRSS) anomaly score, retaining only candidates with N-WRSS $\geq 0.35$ (arbitrary threshold based on figure 21), resulting in a filtered set of 620 planet-like light curves for morphological clustering.
As an initial clustering approach, we employed DBSCAN using a distance matrix derived from Dynamic Time Warping (DTW) Salvador & Chan (2007), a technique well-suited for comparing time-series data with temporal misalignments or non-uniform sampling - conditions typical of TESS light curves.
@@
The results of these clustering experiments are presented thereafter.
4.9.1 Dynamic Time Warping (DTW) and DBSCAN
For the 620 planet-like candidates, we first applied phase folding using the peak period from each curve’s Lomb-Scargle periodogram. The minimum power threshold filtered out the curves that did not have a Lomb Scargle period frequency with power greater than the threshold. The folding was performed using the following function from our pipeline:
clara2.phase_fold(fits_fpath,
                  min_period=0.2,
                  max_period=20,
                  ls_power_thresh=0.05)
This function returns the following key outputs:
  1. 1.
    Phase: Folded phase values of the light curve, sorted, ranging from $[0,1]$.
  2. 2.
    Folded flux: Flux values corresponding to each phase value.
  3. 3.
    Bin centers: Centers of the uniform phase bins used for averaging (default: 100 bins).
  4. 4.
    Bin means: Mean flux values per bin, representing the phase-folded light curve morphology.
To represent each curve’s morphology independently of sampling density, we used the bin means as fixed-length feature vectors. These vectors were padded or truncated to a standard length, normalized, and stored in a list called features. We then computed the Dynamic Time Warping (DTW) pairwise distance matrix using the dtaidistance library:
n = len(features)
D = np.zeros((n, n))
# Compute pairwise DTW distances
for i in range(n):
    for j in range(i+1, n):
        d = dtw.distance_fast(features[i], features[j])
        D[i, j] = D[j, i] = d
With the symmetric distance matrix $D$ computed, we applied DBSCAN for clustering. DBSCAN has two primary hyperparameters:
  1. 1.
    eps: Maximum distance between two points to be considered neighbors.
  2. 2.
    min_samples: Minimum number of neighbors required to form a core point.
To identify optimal clustering parameters, we performed a grid search over combinations of eps and min_samples, evaluating each using the silhouette score. The results are shown in Figure 29.
The two best configurations from the grid search were:
  1. 1.
    $\epsilon=2.0$, min_samples = 10 → 2 clusters, 203 noise points
  2. 2.
    $\epsilon=4.5$, min_samples = 2 → 2 clusters, 29 noise points
Since the DTW-based clustering step is used as a pre-filtering mechanism for removing outliers prior to more refined morphological clustering, we selected the second configuration ($\epsilon=4.5$, min_samples = 2), which yielded a cleaner separation with fewer noise points.
This configuration produced:
  • 29 curves in the DBSCAN noise cluster (1-1),
  • 453 curves in cluster 0,
  • 2 curves in cluster 1.
The two curves in cluster 1 are presented in Figures 27 and 28. Cluster 0, which contained the 453 curves (filtered down from the initial 620 curves), exhibited three distinct morphological types, shown in Figure 30. This cluster also contained 19 TOIs. Neither cluster 1-1 nor cluster $1$ contained any TOIs.
@@ 4.9.2 Feature of a Curve - II
    $s$ $=\begin{cases}\tilde{F}_{i_{\min}-1}-2\tilde{F}_{i_{\min}}+\tilde{F}_{i_{\min}+1},&\text{if }2<i_{\min}<N-3\\ 0,&\text{otherwise}\end{cases}$ (33)
    Simpler explanation: Measures how "pointed" or sharp the dip is at its lowest point.
@@
    $a$ $=\max\left\{\text{ACF}_{k}\mid k>0\text{ and }\text{ACF}_{k}\text{ is a local maximum}\right\}$ (37)
@@
    $\mathcal{T}$ $=\left\{\omega_{i}\mid P(\omega_{i})\text{ is a peak with prominence }>\delta\right\}$ (41)
@@
    We use the Box Least Squares (BLS) algorithm to identify periodic, box-shaped dips in normalized light curves—an effective method for detecting transiting exoplanets. BLS iteratively searches over trial periods to find the one that maximizes a transit signal detection statistic, fitting a simple box model to approximate the flux drop caused by a planet transit. From the best-fit period $P^{\ast}$, we derive the last three of the ten interpretable metrics that characterize the transit-like signal:

### v3

@@ 2.11 Weighted Root Sum Score: A revision of URF anomaly scores
Ensemble fusion provides a natural way to balance the complementary behaviors of URF models with different $\alpha$ values, combining conservative precision and aggressive recall. In particular, the $\alpha=0.3$ and $\alpha=0.9$ models span a broad regime, from conservative anomaly flagging to strong recall, making their intersection an informative testbed for robust score combination.
@@
To give higher preference to better scoring, we set the weights as $w_{3}=0.8$ and $w_{9}=0.2$.
For a generalization of the N-WRSS, let there be $M$ anomaly detection models indexed by $i\in\{1,2,\dots,M\}$, each producing a raw anomaly score $s_{i}^{\prime}$ for a given light curve. We define the generalized weighted root sum score (WRSS) for the intersection set of anomalies flagged by these models as:
$s_{w}=\sum_{i=1}^{M}w_{i}\,\sqrt{s_{i}^{\prime}}$ (17)
where $w_{i}\in[0,1]$ are the weights assigned to each model such that:
$\sum_{i=1}^{M}w_{i}=1$ (18)
The choice of $w_{i}$ allows control over the trade-off between aggressive and conservative anomaly flagging by emphasizing models with desired properties.
For standardized comparison across runs on different test sets, we define the generalized normalized WRSS (N-WRSS) as:
@@
where $\min(s_{w})$ and $\max(s_{w})$ are computed over all curves in the considered set.
@@
The normalized weighted root sum score (N-WRSS $\hat{s}_{w}$) provides a robust ensemble score fusion, capturing the degree of agreement between different URF models ($\alpha_{0.3}$ and $\alpha_{0.9}$ in this case). By computing $\hat{s}_{w}$ over the intersection set — i.e., light curves flagged by both models — we:
@@
  • Mitigate single-model dominance: Linear combinations of normalized scores can be dominated by extreme values from a single model, particularly when models differ in sensitivity or training noise. Applying a monotone concave transformation (e.g., square root) dampens extreme scores, preserving rank order while ensuring the N-WRSS reflects consensus rather than outlier-driven influence. Mathematical intuition: For a monotone concave transformation $f$, we have
    $f(x)-f(y)\leq f^{\prime}(y)(x-y),$ (20)
    with
    $f^{\prime}(y)<1$ (21)
    This shows that large differences between extreme scores are compressed, preserving rank order while reducing outlier influence.
  • Support threshold stability: Dampening extremes enables more robust fine-tuning of N-WRSS thresholds across different datasets, improving reproducibility of anomaly detection.
The empirical benefits of this formulation, in contrast to a simple linear weighted sum score, are demonstrated through Monte Carlo injection-recovery validation in Section 3.6 and Figure 15, where N-WRSS is shown to sustain higher recall and more stable precision-recall tradeoffs across thresholds.
Thus, $\hat{s}_{w}$ on the $\alpha_{0.3}$ - $\alpha_{0.9}$ intersection forms a critical bridge for robust ensemble scoring, providing a stable consensus measure before downstream similarity analysis or labeling. A fuller theoritical treatment of the concavity framework and its implication for ensemble justification is deferred to Section $4.$
@@ 2.12 Astrometric and Stellar Parameter Augmentation
To enable physical interpretation of anomalies detected by URF-4, we enrich high-scoring candidates with astrophysical and astrometric metadata. Specifically, we extract relevant features from the TESS Input Catalog (TIC) and Gaia DR3 for all candidates which exceeded a chosen anomaly threshold (in our case $N-WRSS\geq 0.35$).
@@
Refer to Appendix C for a descriptive list of these features.
@@ 2.13 Evaluating Astrometric Feature Discriminability
To assess which combinations of astrometric and stellar parameters most effectively distinguish high-anomaly candidates, we perform a t-SNE embedding on Sector 1 URF-4 $\alpha_{0.3}\cap\alpha_{0.9}$ set.
@@
This process gives an intuitive insight into discriminative physical features for downstream interpretation, linking anomaly detection outcomes to astrophysical structure.
This method provides an initial quantitative basis for linking astrometric and stellar parameters to anomaly clustering behavior. Future work (4.4) will expand this into a formal statistical validation framework, including comparison with established methods (e.g. D. A. Crake & J. R. Martínez-Galarza (2023); E. J. Melton et al. (2024)) and deeper analysis of why certain features like Teff, RUWE, and $v_{\mathrm{tan}}$ exhibit strong correlations with high N-WRSS anomalies, thereby enabling more grounded astrophysical interpretation of detected outliers.
The results of stellar parameter based t-SNE clustering with N-WRSS overlay showing the correlation between URF scoring and astrophysical features have been demonstrated and discussed in Section $3.$
@@ 2.14 SIMBAD Weak Supervision and Morphological Similarity Matching
To bridge anomaly detection with astrophysical classification, we introduce a weakly supervised labeling approach using the SIMBAD object taxonomy. A subset of high-anomaly light curves (TOIs from Sector 2 - specific choice of sector in this case is arbitrary - for which Simbad Object Labels could be found) is cross-matched with MAST astronomical coordinates to retrieve SIMBAD labels such as Pl, EB*, V*, and others. This offers valuable physical priors without introducing strong biases. These labeled instances are used to construct a Sector 2 reference set of TOIs with known SIMBAD classifications, enabling downstream morphological matching and interpretation. Cosine similarity in the learned feature space propagates these labels to the larger unlabeled set, in line with graph-based semi-supervised learning paradigms (J. E. van Engelen & H. H. Hoos (2020)).
We extract feature vectors from these labeled TOIs using Lomb-Scargle periodograms and flux values of the curves as described in 2.2, apply PCA H. Hotelling (1933) for dimensionality reduction, and save the compressed representations. These features form a labeled reference basis for morphological similarity-based classification. E. Seo et al. (2023)
Subsequently, we compute pairwise cosine similarity between urf-4 $\alpha_{0.3}\cap\alpha_{0.9}$ set candidates and this labeled reference set. Each anomaly is assigned the closest-matching SIMBAD label group based on morphological similarity of the light curve. Following label groups have been defined for this purpose:
@@ 2.14.2 Morphological Classification via Cosine Similarity
    Extracting feature vectors from the input light curve using the same pipeline used for the labeled TOIs as discussed in Section $2.$
@@
Results of cosine similarity based curve labeling have been discussed in Section 3.4, demonstrating, among other things, high confidence morphological similarity based classification of high N-WRSS curves into one of the labels.
@@ 2.15 Clustering and Morphological Filtering Methods
To identify transit-like morphologies, we selected the planet_like classified group of curves from TESS Sectors 1 through 5. We applied a threshold on the normalized weighted root sum square (N-WRSS) anomaly score, retaining only candidates with N-WRSS $\geq 0.35$ (arbitrary threshold based on figure 10 and Section 3.5), resulting in a filtered set of 620 planet-like light curves for morphological clustering.
As a pre-filtering mechanism for clustering, we employed DBSCAN using a distance matrix derived from Dynamic Time Warping (DTW) S. Salvador & P. Chan (2007), a technique well-suited for comparing time-series data with temporal misalignments or non-uniform sampling - conditions typical of TESS light curves. We clustered to have one large cluster with three types of morphologies as presented in Figure 4.
@@ 2.15.1 Feature of a Curve - II
    $s$ $=\begin{cases}\tilde{F}_{i_{\min}-1}-2\tilde{F}_{i_{\min}}\\ \quad+\tilde{F}_{i_{\min}+1},&\text{if }2<i_{\min}<N-3\\ 0,&\text{otherwise}\end{cases}$ (34)
    Simpler explanation: Measures how ”pointed” or sharp the dip is at its lowest point.
@@
[…]
    $\quad\left.\text{ACF}_{k}\text{ is a local maximum}\right\}$ (39)
@@
    $\mathcal{T}$ $=\left\{\omega_{i}\mid P(\omega_{i})\text{ is a peak}\right$. (43)
    $\quad\left.\text{with prominence }>\delta\right\}$ (44)
@@
    We use the Box Least Squares (BLS G. Kovács et al. (2002)) algorithm to identify periodic, box-shaped dips in normalized light curves—an effective method for detecting transiting exoplanets. BLS iteratively searches over trial periods to find the one that maximizes a transit signal detection statistic, fitting a simple box model to approximate the flux drop caused by a planet transit. From the best-fit period $P^{\ast}$, we derive the last three of the ten interpretable metrics that characterize the transit-like signal:
@@
The results and interpretation of clustering based on this new feature set has been demonstrated and discussed in Section 3.4.2, showing how probabilistic clustering on these method yields optimal results, as also justified empirically in Section 3.6.
2.16 Monte Carlo Injection-Recovery Validation
A central challenge in evaluating unsupervised detection frameworks is the absence of ground-truth signals in the majority of survey light curves. Although cross-matching with TOIs provides partial validation, it is incomplete and biased toward previously identified signals. Injection-recovery testing is an established approach (e.g. J. L. Christiansen et al. (2013), J. L. Christiansen et al. (2020)) for quantifying the completeness and reliability of exoplanet detection pipelines. To obtain an absolute measure of sensitivity and to calibrate threshold choices, we performed Monte Carlo injection-recovery experiments using synthetic transits embedded in real TESS light curves. This approach allows us to:
  1. 1.
    Quantify recall and vetting load under controlled conditions.
  2. 2.
    Empirically justify use of NWRSS for quantification of model ensemble consensus.
  3. 3.
    Empirically justify probabilistic methods like GMM for morphological clustering.
  4. 4.
    analyze the false positives that survive URF selection at various thresholds.
2.16.1 Experimental Setup
We constructed a pool of 200 synthetic light curves containing transit signatures spanning a range of depths, noise levels, and orbital periods. For each run, 20 curves were randomly selected from this pool and injected into real TESS light curves. In parallel, 1,980 real SPOC light curves were sampled per run, drawn from Sectors 6–10 and 20–24 and 30, yielding a total of 2,000 curves per run. Random sampling and URF scoring were repeated 80 times to ensure statistical robustness. Both the $\alpha=0.3$ and $\alpha=0.9$ variants were applied to each run, enabling comparison of recovery performance and false-positive behavior across distinct scoring regimes. The full list of the 10 synthetic transit configurations used for injection, including depth and noise ranges, is provided in Table 18. For a single Monte Carlo run, a single random configuration was picked from the ten and then 20 curves of that particular configuration were injected into the test set. A detailed evaluation of recovery performance including confidence intervals across 80 Monte Carlo trials is presented in Sections 3.5 and 3.6.
3 Results
3.1 Correlating Synthetic Parameters to Performance Metrics
Here, we present the systematic trends we observed for these 36 urf-4 sub-variant models. Figures 18, and 19 present the correlation between individual synthetic parameters and the performance metrics TOI Recall, Anomaly Rate and TOI Importance AUC (B.4).
To quantify the relative influence of synthetic training parameters on URF-4 performance, we employed Random Forest regression to model the relationship between our four key synthetic parameters (n_transits, duration_hr (B.3  [^1] During the course of this research, the duration of transit-like events is consistently measured in days across all computations, analyses, and visualizations.), cadence_min, noise_ppm) and three primary evaluation metrics (TOI Recall, Anomaly Rate and TOI Importance AUC). This approach provides interpretable feature importance scores that reveal which synthetic design choices most strongly influence model behavior.
The Random Forest regression analysis, used to quantify how synthetic training parameters influenced URF-4 performance (Figure 5, 6), directly informed the finer-grained experiments presented in Section 2.10.
Figure 5: Feature importance comparison for URF-4 synthetic training parameters. Left panel shows importance scores for TOI Recall prediction, right panel for Anomaly Rate prediction. Transit duration (days B.3) emerges as the dominant factor for both metrics, followed by transit multiplicity. Random Forest regressors trained on 36-variant parameter sweep results with 42-seed random state for reproducibility.
Figure 7 presents the performance trajectories of the top 10 URF-4 variants ranked by TOI Recall AUC across varying anomaly threshold percentiles. It shows the top URF-4 optimal models achieving  40% recall in the top 25% of anomaly scores. However, while individual feature importance rankings guide parameter prioritization for focused optimization (2.9), the combined metric analysis 2.7 becomes essential for interpreting current results since different parameters optimize different performance aspects. The divergent trajectories between TOI Recall AUC and TOI Importance AUC in Figure 7 highlight the need for weighted combination scoring to balance detection completeness against astrophysical relevance in URF-4 variant selection.
3.2 Results: Generalization Evaluation of Balanced URF Variant ($\alpha$=0.5)
Evaluation across 10 random subsets ($\approx$40,000 total light curves) yields Table 2. TOI Recall AUC deviates only 0.014 below target, indicating consistent TOI prioritization across the anomaly score distribution. TOI Importance AUC exceeds the target by 0.023, suggesting the model performs even better than expected at detecting astrophysically significant targets. The binary recall of 32.1% and anomaly rate of 27.9% indicate closely related TOI Recall and Anomaly Rate as predicted by Figure 18. The value of these metrics will be improved by predicting over a finer mesh of synthetic design features and using a more comprehensive model scoring metric as presented in Section 2.10.
Metric Value Target Deviation
Mean TOI Recall AUC 0.547 0.561 -0.014
Mean TOI Importance AUC 0.143 0.120 +0.023
Mean Binary TOI Recall 0.321 - -
Mean Anomaly Rate 0.279 - -
Verdict: This model satisfies our generalization criteria (that it will perform consistently on similar data), with both metrics falling within the acceptable ±0.03-0.05 deviation range.
Metric $\alpha$ = 0.3 $\alpha$ = 0.5 $\alpha$ = 0.9 Best Performer
TOI Recall AUC 0.521 0.547 0.557 $\alpha$ = 0.9
TOI Importance AUC 0.141 0.143 0.139 $\alpha$ = 0.5
Mean TOI Importance (Top 20%) 0.393 0.422 0.359 $\alpha$ = 0.5
Binary TOI Recall 0.493 0.321 0.682 $\alpha$ = 0.9
Anomaly Rate 0.447 0.279 0.558 $\alpha$ = 0.5
Metric S1 Best S2 Best S3 Best S4 Best S5 Best Consistency
3.2.1 Results: Alpha Variant Performance Comparison
3.2.2 Intepretation of Results of Table 3
  1. 1.
    $\alpha$-parameter control works as designed
  2. 2.
    Behavioral differences are robust
  3. 3.
    Trade-offs are quantifiable
3.2.3 Results: Cross-Sector Generalization of Alpha Variants
Operational Implications: The predictable cross-sector behavior enables confident deployment of $\alpha$ variants to new TESS sectors without retraining.
3.3 Results: t-SNE Clustering on Stellar Parameters with N-WRSS overlay
The most visually separable and physically interpretable feature set was: [’Teff’, ’vtan’, ’ruwe’], showing excellent concentrated red region (high N-WRSS) clusters. These three features showed up consistently across higher silhouette rankings of feature combinations. Testing it on sector 1 data showed the t-SNE clustering as visualized in Figure 8
Figure 8: t-SNE clustering of Sector 1 N-WRSS $\geq$ 0.6 curves based on [Teff, vtan, ruwe]. N-WRSS goes from blue to red increasing from 0.0 to 0.9 (non-linearly to mark N-WRSS $\geq$ 0.6 better visually)
We performed cross sector t-SNE tests across TESS SPOC Sectors 1-4 based on the feature set $T_{eff}$, $v_{tan}$, RUWE]. The t-SNE embeddings in Figure 9 reveal consistent structural patterns across TESS Sectors 1–4.
The projection shows a recurring concentration of high N-WRSS scores (red regions) along narrow manifolds in the reduced space, suggesting the existence of a physically distinct anomaly subpopulation. This reinforces the hypothesis that high-scoring anomalies are not randomly distributed but instead correlate with specific stellar or kinematic regimes, validating the physical relevance of the URF-derived anomaly scores across sectors.
3.4 Results of Morphological Classification via Cosine Similarity
To assess the astrophysical plausibility of morphologically matched labels, we analyze the distribution of URF anomaly scores across the assigned SIMBAD (M. Wenger et al. (2000a)) label categories.
  1. 1.
  2. 2.
  3. 3.
3.4.1 Interpretation of Cosine Similarity Results
The three diagnostic views (figures 10, 11 and 12) enable a structured assessment of how anomaly scores and morphological similarity align with known astrophysical classes.
Second, the high-similarity/low-anomaly set (Figure 11) demonstrates that URF correctly assigns low anomaly scores to morphologically standard light curves which can be classified into a labeled category. This supports the argument that low URF scores represent “normal” behavior in the training distribution.